# AxLSTMs: learning self-supervised audio representations with xLSTMs

### v3

@@ 2.1 Prerequisites: xLSTM and the mLSTM block
${C_{t}}$ $=\ {f_{t}}\ {C_{t-1}}\ +\ {i_{t}}\ {v_{t}\ k_{t}^{\top}},$ (1)
${n_{t}}$ $=\ {f_{t}}\ {n_{t-1}}\ +\ {i_{t}}\ {k_{t}},$ (2)
$h_{t}$ $=\ {o_{t}}\ \odot\ {C_{t}}{q_{t}}\ /\ \max\left\{|{n_{t}^{\top}}{q_{t}}|,1\right\},$ (3)
$q_{t}$ $=\ W_{q}\ x_{t}\ +\ b_{q}$ (4)
$k_{t}$ $=\ \frac{1}{\sqrt{d}}W_{k}\ x_{t}\ +\ b_{k}$ (5)
$v_{t}$ $=\ W_{v}\ x_{t}\ +\ b_{v}$ (6)
${i_{t}}$ $=\ \exp{\left(\ w^{\top}_{i}\ x_{t}\ +\ b_{i}\right)}\ ,$ (7)
${f_{t}}$ $=\ \exp{\left(\ w^{\top}_{f}\ x_{t}\ +\ b_{f}\right)},$ (8)
$o_{t}$ $=\ \sigma\left(\ W_{o}\ x_{t}\ +\ b_{o}\right)\,$ (9)
@@ 2.2 AxLSTM: Modelling masked patches with xLSTMs
Table 2: Comparing AxLSTMs with popular self-supervised audio representations. We used pretrained models from cited papers to extract fixed feature vectors and conducted our own downstream experiments. LS, AS, VP, LL stand for LibriSpeech, AudioSet, VoxPopuli and LibriLight datasets, respectively. Original SSAST [6] was trained on AS+LS, whereas we pretrained the directly comparable underlined SSAST baselines (SSAST-T,S,B). *includes decoder parameters
@@ 3.2 Implementation details
Pretraining: All the proposed models are pretrained on randomly cropped 2-second audio clips, resulting in an input spectrogram of shape $[\mathtt{200}\times\mathtt{80}]$. Our default configuration consists of $l=12$ number of stacked mLSTM blocks with a model feature dimension of $d_{m}=192$, the same as those of a ViT-Tiny [20] encoder, but we also evaluate Small ($d_{m}=$384, $l$=12) and Base ($d_{m}$=768, $l$=12) encoder configurations. By default, a patch embedding layer that computes $[4\times 16]$ shaped non-overlapping patches (along time and frequency, respectively) is used in all our experiments. While [16, 17] suggest a default expansion factor of $E_{f}=2$, in early ablations we found that $E_{f}=3$ performed the best (Section 4.2). Ablation experiments also revealed that disabling the flip operation performed better than enabling flip in alternating blocks. We build on hyperparameter recommendations from previous work [6, 14]: all models are trained for 100 epochs with a batch size of 1024 and a weight decay of 0.05. AdamW optimizer with a linear warmup for 10 epochs followed by a cosine learning rate decay schedule is used. No data augmentations were used.
@@ 4.1 Comparison with existing works
Table 2 shows how the proposed AxLSTM models fare against recent audio representations. Directly comparable SSAST and AxLSTM feature representations have identical feature embedding sizes, however, feature vector sizes extracted using other referred methods can vary. While sub-optimal, it is infeasible to retrain all pretrained representations to have the same embedding sizes, and our evaluation protocol is in line with recent frameworks for evaluating self-supervised audio representations [23, 34]. SSAST [6] represents the officially released model trained on AudioSet+LibriSpeech, whereas underlined SSAST representations were pretrained by us. We can see that AxLSTM models consistently outperform their transformer-based SSAST counterparts by a considerable margin, while having over 23% fewer parameters, with AxLSTM-B yielding over 25% relative improvement in aggregate performance (86.6±0.2 v/s 68.7±0.3). Despite being based on the older SSAST framework, our AxLSTM-B model matches the more recent masked modeling-based approaches such as AudioMAE [7], BEATS-iter3 [19] and MSM-MAE-208 [21] in overall performance, while having over 25% fewer parameters. The recent Mamba-based SSAM [14] models outperform comparable AxLSTM models in overall performance, however, it’s worth noting that AxLSTM models outperform SSAM models in music and pitch perception tasks (BO, Mri-T, NSynth-5h) whereas SSAM models shine in Speech-based and Audio classification tasks, performing notably better for emotion recognition (CD) and spoken language identification (VL). Finally, the recent MWMAE-Base [22] model with local-global self-attention outperforms every evaluated approach, including AxLSTM-B, albeit with notably more parameters. Overall, while the AxLSTM model doesn’t establish a new state-of-the-art, it significantly outperforms SSAST-based transformer baselines while matching the performance of several other newer masked modeling approaches while possessing fewer parameters, thus demonstrating the viability of xLSTMs to learn self-supervised audio representations.

### v4

@@ 2.1 Prerequisites: xLSTM and the mLSTM block
${C_{t}}\$$=\ {f_{t}}\ {C_{t-1}}\ +\ {i_{t}}\ {v_{t}\ k_{t}^{\top}},$ (1)
${n_{t}}\$$=\ {f_{t}}\ {n_{t-1}}\ +\ {i_{t}}\ {k_{t}},$ (2)
$h_{t}\$$=\ {o_{t}}\ \odot\ {C_{t}}{q_{t}}\ /\ \max\left\{|{n_{t}^{\top}}{q_{t}}|,1\right\},$ (3)
$q_{t}\$$=\ W_{q}\ x_{t}\ +\ b_{q}$ (4)
$k_{t}\$$=\ \frac{1}{\sqrt{d}}W_{k}\ x_{t}\ +\ b_{k}$ (5)
$v_{t}\$$=\ W_{v}\ x_{t}\ +\ b_{v}$ (6)
${i_{t}}\$$=\ \exp{\left(\ w^{\top}_{i}\ x_{t}\ +\ b_{i}\right)}\ ,$ (7)
${f_{t}}\$$=\ \exp{\left(\ w^{\top}_{f}\ x_{t}\ +\ b_{f}\right)},$ (8)
$o_{t}\$$=\ \sigma\left(\ W_{o}\ x_{t}\ +\ b_{o}\right)\,$ (9)
@@ 2.2 AxLSTM: Modelling masked patches with xLSTMs
Table 2: Comparing AxLSTMs with popular self-supervised audio representations. We used pretrained models from cited papers to extract fixed feature vectors and conducted our own downstream experiments. LS, AS, VP, $L$ stand for LibriSpeech, AudioSet, VoxPopuli and LibriLight datasets, respectively. Original SSAST [6] was trained on AS+LS, whereas we pretrained the directly comparable underlined SSAST baselines (SSAST-T,S,B). *includes decoder parameters
@@ 3.2 Implementation details
Pretraining: All the proposed models are pretrained on randomly cropped 2-second audio clips, resulting in an input spectrogram of shape $\mathtt{200}\times\mathtt{80}]$. Our default configuration consists of $l=12$ number of stacked mLSTM blocks with a model feature dimension of $d_{m}=192$, the same as those of a ViT-Tiny [20] encoder, but we also evaluate Small ($d_{m}=$384, ll=12) and Base ($d_{m}$=768, ll=12) encoder configurations. By default, a patch embedding layer that computes $4\times 16]$ shaped non-overlapping patches (along time and frequency, respectively) is used in all our experiments. While [16, 17] suggest a default expansion factor of $E_{f}=2$, in early ablations we found that $E_{f}=3$ performed the best (Section 4.2). Ablation experiments also revealed that disabling the flip operation performed better than enabling flip in alternating blocks. We build on hyperparameter recommendations from previous work [6, 14]: all models are trained for 100 epochs with a batch size of 1024 and a weight decay of 0.05. AdamW optimizer with a linear warmup for 10 epochs followed by a cosine learning rate decay schedule is used. No data augmentations were used.
@@ 4.1 Comparison with existing works
Table 2 shows how the proposed AxLSTM models fare against recent audio representations. Directly comparable SSAST and AxLSTM feature representations have identical feature embedding sizes, however, feature vector sizes extracted using other referred methods can vary. While sub-optimal, it is infeasible to retrain all pretrained representations to have the same embedding sizes, and our evaluation protocol is in line with recent frameworks for evaluating self-supervised audio representations [23, 34]. SSAST [6] represents the officially released model trained on AudioSet+LibriSpeech, whereas underlined SSAST representations were pretrained by us. We can see that AxLSTM models consistently outperform their transformer-based SSAST counterparts by a considerable margin, while having over 23% fewer parameters, with AxLSTM-B yielding over 25% relative improvement in aggregate performance (86.6±0.2 v/s 68.7±0.3). Despite being based on the older SSAST framework, our AxLSTM-B model matches the more recent masked modeling-based approaches such as AudioMAE [7], BEATS-iter3 [19] and MSM-MAE-208 [21] in overall performance, while having over 25% fewer parameters. The recent Mamba-based SSAM [14] models outperform comparable AxLSTM models in overall performance, however, it’s worth noting that AxLSTM models outperform SSAM models in music and pitch perception tasks (BO, Mri-T, NSynth-5h) whereas SSAM models shine in Speech-based and Audio classification tasks, performing notably better for emotion recognition (CD) and spoken language identification (VL). Finally, the recent MWMAE-Base [22] model with local-global self-attention outperforms every evaluated approach, including AxLSTM-B, albeit with notably more parameters. Overall, while the AxLSTM model doesn't establish a new state-of-the-art, it significantly outperforms SSAST-based transformer baselines while matching the performance of several other newer masked modeling approaches while possessing fewer parameters, thus demonstrating the viability of xLSTMs to learn self-supervised audio representations.